### v3

# Large Language Models Develop Novel Social Biases Through Adaptive Exploration

Abstract
As large language models (LLMs) are adopted into frameworks that grant them the capacity to make real decisions, it is increasingly important to ensure that they are unbiased. In this paper, we argue that the predominant approach of simply removing existing biases from models is not enough. Using a paradigm from the psychology literature, we demonstrate that LLMs can spontaneously develop novel social biases about artificial demographic groups even when no inherent differences exist. These biases result in highly stratified task allocations, which are less fair than assignments by human participants and are exacerbated by newer and larger models. In social science, emergent biases like these have been shown to result from exploration-exploitation trade-offs, where the decision-maker explores too little, allowing early observations to strongly influence impressions about entire demographic groups. To alleviate this effect, we examine a series of interventions targeting model inputs, problem structure, and explicit steering. We find that explicitly incentivizing exploration most robustly reduces stratification, highlighting the need for better multifaceted objectives to mitigate bias. These results reveal that LLMs are not merely passive mirrors of human social biases, but can actively create new ones from experience, raising urgent questions about how these systems will shape societies over time. 00footnotetext: Our code is available at github.com/addisonwu05/LLM-Emergent-Stereotypes.
1 Introduction
As LLMs become embedded in everyday applications across countless tasks, it is imperative for them to be unbiased, meaning that they treat people equally across racial, gender, and other social groups. This is critical because biased behavior in such systems can perpetuate and amplify existing societal inequities, undermine user trust, and lead to systematically unequal access to resources and opportunities. However, current LLMs are biased: they mirror existing human biases (e.g., Bolukbasi et al., 2016; Caliskan et al., 2017; Dhamala et al., 2021; Nadeem et al., 2021; Tamkin et al., 2023), and many efforts have been dedicated towards removing these biases (e.g., Bordia and Bowman, 2019; Guo et al., 2022; Liang et al., 2021; Meade et al., 2022; Yu et al., 2023). This process has proven to be challenging, as models that pass benchmarks continue to reveal subtle discriminatory behaviors (Bai et al., 2025b; Hofmann et al., 2024; Ji et al., 2025; Zipperling et al., 2025).
In this paper, we argue that removing existing biases is only one aspect of the problem. Like people, LLMs can also invent novel biases that influence human and agent behavior. Stereotype biases in humans can naturally emerge through experiences that constrain exploration (Bai et al., 2022a; 2025a; Fang and Moro, 2011; Merton, 1948; Schelling, 1971): residents search only familiar neighborhoods, reinforcing segregation (Krysan and Crowder, 2017); police repeatedly patrol high-crime areas, disproportionately arresting minorities (Lum and Isaac, 2016); managers avoid hiring unconventional candidates, maintaining incorrect beliefs (Baek and Makhdoumi, 2023); and individuals view a group negatively after one bad encounter, escalating conflicts (Denrell and March, 2001). This mechanism parallels the exploration-exploitation dilemma in reinforcement learning (Ensign et al., 2018; Sutton et al., 1998): when iteratively facing choices with multiple options, each choice is costly but informative, forcing decision-makers to balance exploring novel options with exploiting what worked before. This phenomena becomes pertinent at a time when foundation models are being integrated into agentic frameworks, letting them retain persistent belief states across interactions, while also granting them autonomy to make decisions with limited human oversight (Krishnamurthy et al., 2024; Laskin et al., 2023; Raparthy et al., 2024; Shinn et al., 2023).
We illustrate this process of developing novel biases using a hiring game paradigm from the psychology literature (Bai et al., 2022a; 2025a). Participants act as hiring managers to allocate a series of jobs, each of which has candidates from four artificial demographic groups, and they are rewarded for how many hired candidates succeed. Jobs are split into four types along two psychological dimensions, warmth and competence (Fiske et al., 2002), following human data from Bai et al. (2025). For example, doctors are seen as trustworthy and competent while janitors are viewed as less so (Koenig and Eagly, 2014). Unknown to the participant, all candidates are equally likely to succeed with probability at each job. However, as participants explore by assigning candidates to roles and receive feedback on whether they succeed, these early observations often lead them to form inaccurate impressions about the underlying traits of each group, leading them to stratify candidates by assigning different groups to different job types. In other words, people do not explore enough to remove biases caused by inherently random feedback, causing them to treat groups unequally despite no real differences. Afterwards, people retained these biases, rating certain groups as more competent or caring than others. This process demonstrates how humans can develop new biases simply from engaging in sequential decision-making with noisy outcomes.
When LLM decision-makers are put in similar situations, do they also develop novel biases from insufficient exploration? We test this by replicating the iterative hiring experiment on LLMs (Figure 1), prompting them to complete it using multi-turn dialogue (Section 3). Our results demonstrate that not only do LLMs develop new biases, but LLMs also assign different jobs to demographic groups with even more stratification than human participants. Furthermore, newer and larger models show increased stratification effects, suggesting a dangerous trend that models with higher reasoning capabilities lead to more unequal outcomes (Section 4). In follow-up experiments, we investigate the generality of our findings using two other multi-turn decision settings, along with a series of bias mitigation interventions focused on increasing exploration (Section 5). Compared to other strategies, explicitly incorporating diversity in the prompted objective is most effective for reducing stratification behaviors in LLMs. This result illustrates the importance of defining multifaceted goals that incorporate societal values when instructing modern AI systems, allowing us to leverage these powerful instruction-followers toward socially desirable outcomes.
Our findings reflect a general, recurring theme in optimization and AI — that stronger optimizers require better-formulated goals (Amodei et al., 2016; Hadfield-Menell et al., 2017; Manheim and Garrabrant, 2018; Pan et al., 2022; Smith and Winkler, 2006). As a concrete example, consider the contrast between newspapers and social media, which share the objective of increasing audience engagement. While newspapers were limited by lack of feedback, social media platforms used closed-loop optimization with user data to improve recommendations—but this led to negative societal consequences such as echo chambers and polarization (Allcott et al., 2020; Bakshy et al., 2015; Cinelli et al., 2021). Our results show that LLMs as optimizers have also outgrown simple reasoning objectives. To adapt to the improved capabilities that state-of-the-art models provide, we believe that holistic objectives that incorporate societal values (Bai et al., 2022c; Klingefjord et al., 2024) are imperative to ensure that AI systems stay unbiased as they explore and interact with the world.
2 Related work
2.1 Quantifying and addressing biases in LLMs
Stereotype biases in language models are well recognized as a long-standing problem, from word embeddings (Bolukbasi et al., 2016; Caliskan et al., 2017) to autoregressive models (Dhamala et al., 2021; Liang et al., 2023; Nadeem et al., 2021; Huang et al., 2025). To evaluate these biases, benchmarks have mainly focused on existing categories embedded in society, such as race (Hofmann et al., 2024; Wang et al., 2023), gender and sexual orientation (Ovalle et al., 2023; Wan et al., 2023), age (Tamkin et al., 2023), religion (Abid et al., 2021), occupation (Kirk et al., 2021), and cultural background (Shen et al., 2024). To reduce these biases, intervention techniques also target known stereotypes by creating alignment datasets (Bai et al., 2022b; Zhang et al., 2025), editing model activations (Prakash and Roy, 2024; Sun et al., 2025; Yu and Ananiadou, 2025), or prompting (Si et al., 2023). While useful for addressing existing biases, these approaches cannot capture or address new forms of bias that emerge as models interact with the world and adapt their beliefs. Here, we show that LLMs can generate entirely novel and potentially problematic biases, unseen in any data.
2.2 Challenges for exploration with LLMs
In-context learning illustrates how LLMs can generalize from very few examples without training, leading to superior performance on many tasks (akyürek2023what; Brown et al., 2020; Shi et al., 2024). However, in this paradigm, LLMs have also displayed notable shortcomings when operating in unfamiliar distributions or on tasks that require generalization beyond surface patterns. For example, in multi-armed bandit tasks, LLMs tend to fixate on the same option that first results in a successful reward, even though this is suboptimal (Krishnamurthy et al., 2024; Pan et al., 2025; Schmied et al., 2025). LLMs can also make spurious and incorrect generalizations from confounded in-context data, prioritizing surface-level features such as sentiment (Fei et al., 2023), length (Schoch and Ji, 2025), or those favored in its priors (Si et al., 2023). More broadly, LLMs display inductive biases toward simpler or more common patterns (McCoy et al., 2024a; b), which can lead to them over-indexing on such patterns within in-context data (Li et al., 2025; Liu et al., 2025). Together, these results highlight how limited exploration—through fixation, spurious correlations, or early lock-in on presumed patterns—remains a central bottleneck to robust generalization.
3 Methodology
3.1 Hiring game paradigm
Imagine being hired as a consultant by the mayor of a fictional city. Your task is to help fill twenty jobs such as doctors, lawyers, childcare aides, janitors with applicants from four unfamiliar demographic groups: Tufa, Aima, Reku, and Weki. In each round, there is a new job vacancy and four applicants, one from each group, awaiting your decision. Once you make your choice, you learn immediately whether the hire was successful, and move on to the next round. Your goal is to maximize successful hires across 40 rounds, which will be converted into a real bonus compensation.
This simple contextual multi-armed bandit setup from Bai et al. (2025) is designed to strip away existing biases: participants belonged to none of the groups—reducing in-group loyalty (Brewer, 1979), clear instructions and short trials minimized cognitive load (Macrae et al., 1994), and job candidates had equal population sizes to prevent data imbalance (Fiedler, 2000). Crucially, unknown to participants, the odds of success were identical for every group and every job. At each round, whether any job is a good fit for any selected applicant is a random variable sampled from .
In the original experiment, human participants failed to realize that there were no meaningful differences among groups. Instead, they became entrenched in their own successes: once they observed that a Tufa was a good doctor or a Weki worked well as a janitor, participants kept repeating similar choices rather than exploring alternatives. In doing so, they inadvertently built a stratified city of their own making, and created new mental stereotypes imagining Tufas as warm and competent while casting Wekis as untrustworthy and incompetent (Bai et al., 2025a). This experiment provides the baseline human data for our evaluation of LLMs (see Appendix C for details), which we test using the same hiring task.
3.2 Metrics
We introduce three complementary metrics to quantify stereotype emergence. The first measure, stratification index (SI), reflects how strongly groups concentrate in specific job classes. The second measure, between-group divergence (BGD), captures whether groups’ assigned job classes diverge from one another. The third metric, group assignment stochasticity index (GASI), assesses whether observed stereotypes are consistent across runs.
Throughout this section, let denote the set of demographic groups, the collection of independent runs of the hiring game, and the set of 4 job classes: high competence and high warmth (e.g., doctor), high competence and low warmth (e.g., lawyer), low competence and high warmth (e.g., childcare aide), and low competence and low warmth (e.g., janitor) (Bai et al., 2025a; Fiske et al., 2002; Fiske and Dupree, 2014; Koenig and Eagly, 2014). For each group in run , we write for its empirical allocation distribution over the job classes, and for the uniform distribution on . and JSD denote entropy and Jensen-Shannon divergence over probability distributions, respectively, with all logarithms calculated using base 2.
Stratification Index (SI)
SI measures how much the decision-maker funnels each demographic into particular classes of jobs, rather than distributing them uniformly across different classes.
When jobs are uniform across , including our experimental settings, SI is also equivalent to the expected mutual information between and across runs (proof in Appendix B.1.1).
Between-Group Divergence (BGD)
If each demographic is funneled into its own subset of jobs, BGD measures how different these group-specific allocation patterns are from one another.
Group Assignment Stochasticity Index (GASI)
One reasonable concern is whether the observed biases are instead reflections of subtle underlying associations (e.g., with artificial demographic names or positional biases). GASI measures how consistently group–role associations recur across independent runs: low stochasticity suggests latent, ingrained biases, whereas high stochasticity means that the observed patterns arise due to emergent dynamics within each run.
Appendix B contains numerical analyses for each metric—showing they capture distinct and complementary aspects of stereotype emergence, and interpretations for each metric’s range of values.
4 Do LLMs naturally segregate equal groups?
4.1 Models and hyperparameters
We examined a variety of state-of-the-art LLMs and their predecessors, both proprietary and open-source: GPT-[3.5, 4o], Claude [3 Haiku, 4 Sonnet], Gemini [1.5, 2.0, 2.5] Flash, Qwen 2.5-[7B, 72B] Instruct Turbo, Llama [3.2 3B, 11B, 90B, 4 Scout 17B-16E, 4 Maverick 17B-128E] (frontier models of each family are in bold). In addition, we tested two reasoning models, one proprietary—OpenAI o3, and one open-source—DeepSeek-R1. Each model was prompted at its default temperature, with both direct and chain-of-thought prompting (CoT; Wei et al., 2022). For reasoning models, the default medium reasoning effort was used. For each model and prompt type, we collected runs of the 40-round hiring game from Section 3.1, with the order of jobs shuffled each run. Prompts are in Appendix A.1.
4.2 Results
Frontier models develop biases and stratify even more severely than humans.
Our experiments find that LLMs develop emergent biases as they explore, with frontier models stratifying groups into different job classes at an even higher degree than people. As depicted in Figure 2, human participants produced stratified allocations (SI = , 95% CI ; BGD = ) far beyond what occurs when conducting fair random assignments (SI = , 95% CI ; BGD = ). However, all frontier LLMs produced even more stratified outcomes than humans (mean SI = , mean BGD = ). Among non-reasoning models, Claude Sonnet 4 with direct prompts stratified the most (SI = 1.79, 95%-CI whereas Qwen 2.5-72B with CoT (SI = 0.89, 95%-CI ) was closest to human levels. Reasoning models also stratified more extremely (OpenAI o3 SI = 1.83, BGD = .80; DeepSeek-R1 SI = 1.41, BGD = .71). Furthermore, we confirmed high stochasticity in group-job assignments (mean GASI = vs. human = , Table 1) across many models and prompts. This suggests that stratification patterns are learned during each run (e.g., through sampled candidate successes), rather than originating from training data (more analyses in Appendix F).
Claude Sonnet 4 Gemini 2.5 Flash DeepSeek–R1 Llama 4 Maverick GPT–4o Qwen 2.5 72B OpenAI o3 Humans Prompt CoT Direct CoT Direct Reasoning CoT Direct CoT Direct CoT Direct Reasoning - GASI 0.61 0.30 0.60 0.60 0.57 0.56 0.52 0.51 0.56 0.50 0.45 0.48 0.47
Newer and larger models have a greater tendency to stratify compared to predecessors.
In experiments across each model family {Claude, GPT, Gemini, Llama3.2, Llama4, Qwen2.5}, we observe that newer and larger models stratified statistically significantly more as measured by both SI and BGD (Figure 3). For instance, Claude 4 Sonnet’s SI was more than eight times Claude 3 Haiku’s in the direct prompting condition. This runs contrary to results on single-prompt bias benchmarks such as BBQ, where newer and larger models consistently demonstrate higher performance than predecessors (Center for Research on Foundation Models, 2025; Liang et al., 2023; Parrish et al., 2022). We provide a visualization of the strong negative correlation between BBQ score and negative SI (adjusted so that higher = better) in Figure 4. We also provide a visualization for how newer models within each family stratify more using single-run rank-ordered job allocation matrices (Appendix D). Our finding that stronger models are more at risk of developing emergent biases highlights the need to specifically attend to this new kind of bias.
5 Interventions to determine factors behind stratification
To understand the sources of LLMs’ stratification and test potential solutions, we performed three types of interventions. First, we varied model-specific inputs such as temperature and CoT prompting, which marginally reduced stratification (Section 5.1). Next, we altered structural features of the task environment, including testing alternate settings and removing gamified rewards given to the LLM—which did not mitigate stratification, and changing success rates and adding more features—which led to reduced stratification, though not robustly (Section 5.2). Finally, we tested a collection of prompt steers focusing on LLMs’ values, community norms, or the explicit objective function in the scenario. Most approaches were partially successful, but explicitly asking the model to optimize for diversity was most robust and effective, showing particular promise as an applicational intervention (Section 5.3).
5.1 System-level interventions
Chain-of-thought prompting does not meaningfully reduce stratification.
CoT has shown promise in encouraging exploration and reducing bias (Gupta et al., 2025; Krishnamurthy et al., 2024), and is a general strategy to improve performance (Wei et al., 2022). While CoT decreased stratification in most frontier models (Figure 2), these changes were often not statistically significant. With CoT, Qwen 2.5 72B—the lowest SI frontier model—reduced stratification to within human ranges. However, all outcomes were still far more stratified than fair random assignments.
Counterintuitively, neither does increasing temperature.
Another standard strategy to encourage randomness is to increase model temperature (Du et al., 2025). We test this by prompting each frontier model (except Claude 4 Sonnet whose maximal temperature is ) with an increased temperature of for runs. We report only direct prompting results, as CoT devolved outputs into gibberish after 7-10 rounds at and for all models. For direct prompts, increasing the temperature to did not produce statistically significant reductions in stratification for Gemini 2.5 (), GPT-4o (), or Llama 4 Maverick (). While we observed a statistically significant decrease in stratification for Qwen 2.5-72B (), the resultant SI of remained above the human baseline—well within the high-stratification regime.
These insufficient interventions aimed at fixing system behaviors suggest that emergent biases in LLMs are not merely a byproduct of poor reasoning or limited sampling diversity, but reflect a deeper structural tendency in their allocation behavior.
5.2 Structural interventions
New decision settings without gamified rewards yield similar stratification effects.
To confirm that stratification is not caused by our specific setup, we test two additional settings with similar multi-turn decisions: refugee resettlement (Bansak et al., 2018; 2016) and military conscript assignment (Sørlie et al., 2020). Starting from the same multi-turn allocation paradigm, we replaced categorized jobs with either geographically-clustered cities in a country or military camps from different divisions. In the resettlement setting, we also replaced the fictional demographics with low-resource indigenous ethnicities from Central Asia for further realism, confirming that initial biases across ethnicities are spurious (across all conditions GASI ). While the original experiment from Bai et al. (2025a) used a points system for successful job assignments to incentivize participants, these incentives are not necessary for LLMs. Thus, our new settings remove the points system and only instruct the LLM to maximize successful assignments. See Appendices A.4 and A.5 for prompts and Appendix E for full results.
In both settings, we observed strong stratification effects. Across five frontier models and direct/CoT prompts, we observed mean SIs of 1.13 and 1.26 for refugee resettlement and conscription assignment, respectively. These results show that the emergent biases generalize across domains, and that they are not dependent on explicit gamified rewards that are only introduced in pseudo-realistic scenarios.
Lowering success probabilities reduces but does not remove stratification.
At first glance, biases developed from exploration may be a result of high success rates, where exploration is not necessary to perform well. To test this hypothesis and widen our scope of problems, we replicated the experiment with reduced success rates of 0.1 for all candidate-job pairs. Due to cost constraints, we excluded reasoning models. As shown in Figure 5, this encouraged more exploration and produced less stratified outcomes, with more pronounced reductions using CoT. Notably, for Llama 4 Maverick, direct prompting resulted in biased allocations (mean SI = 1.23), whereas CoT drastically reduced this tendency (mean SI = 0.31). However, only GPT-4o’s direct assignments and Claude 4 Sonnet’s CoT assignments had SIs below the random threshold, indicating that success rates are not the only factor behind stratification. These tests with lower success rates show that more challenging environments can partially offset formation of premature biases, but at the cost of being artificial—raising the question of how naturalistic difficulties would push models to structure allocations.
Using realistic job-wise success probabilities limits these stratification reductions.
We follow the previous intervention with a variant that assigns job success probabilities equal to the LLM’s elicited prior. Conducted using the fairest model in the setting (GPT-4o), we set success probabilities for each job by asking the LLM what percentage of the general population would succeed in the role. These values ranged from 6–87%, with each of the four job types (high/low warmth high/low competence) following a different distribution. See Appendix A.3 for prompts and job success probabilities. With these new probabilities, GPT-4o’s allocations were no longer close to fair random assignment, with SIs of 0.82 for direct and 0.60 for CoT. While stratification did decrease from the condition, GPT-4o was unable to replicate the ideal levels it attained in the setting, suggesting that LLMs remain likely to stratify in real-world settings.
Models form incorrect biases even when success probabilities per demographic are different.
While previous experiments show how LLMs develop biases when demographics are equal, next we illustrate how this persists even when some demographics perform better than others. We test this by modifying success rates: Each demographic is best at a job category (with success ), worst at a job category (), and moderate ( and ) at the other two. Once the hiring rounds are complete, we additionally ask the LLM to identify the demographic that is most likely to succeed at a job from each job quadrant. More details are in Appendix J.
We tested GPT-4o and Gemini-2.5-Flash for both direct and CoT prompts, in both 40 and 80-round hiring setups with 30 trials per setting. In the 40-round setup, LLMs only identified the best-performing group 26.3% of the time, barely surpassing random chance. It mistakenly identifies the second-best group 29.0% of the time, the third best group 21.5% of the time, and even the worst-fitting group 23.3% of the time (95% CIs shown in Appendix J, Table 4). In the condition where LLMs were given 80 hiring rounds and explicitly informed of this, there was no statistically significant difference in accuracy vs. the 40-round case (26.2%, 30.5%, 24.4%, 18.9%; Appendix J Table 5). Models were so easily influenced by initial feedback that they were unable to adapt their exploration to attain better long-term benefits.
Providing more information about candidates can help reduce stratification.
Another case to consider is scenarios where the LLM has access to richer information beyond group labels alone. Real-world decision making can involve multiple dimensions of context, and incorporating additional features allows us to explore if stratification arises when models can explain observations using other available features. We examined this question using the refugee resettlement setting with established realistic feature from Bansak et al. (2018; 2016): age and education. For experiment details and prompts, see Appendix A.4.
We find that as we add additional features, most models shift progressively towards less stratification by ethnic group (Figure 6(a)). However, the degree of this shift varied by model and prompting method. For example, CoT prompts led to fairer assignments across almost all models and feature combinations. On the other hand, while Claude 4 Sonnet stratified less than other models without new features, adding features did not always make its assignments more fair. Other models generally saw decreases in stratification with more features, with most attaining SIs in proximity to random assignment, but Gemini and Claude retained relatively higher SIs around 0.7. This indicates that while LLMs can explain observed feedback using other available features, they may also still anchor to spurious demographic signals.
However, the type of additional information modulates reductions in stratification.
While we use the most prevalent features (age, education) for the resettlement task in our previous analysis (Bansak et al., 2018; 2016), in real world applications a myriad of features could be available for individuals. Thus, it is imperative to distinguish whether arbitrary features equally increase exploration by expanding the hypothesis space, or if LLMs selectively adjust stratification based on the additional features’ contextual importance. To examine this, we replicate the resettlement experiment using two comparatively less salient features: hair color and tattoo shape (Martin et al., 2014). We observe substantially higher levels of stratification with these features (Figure 6(b)), with mean reductions in SI of 0.25, 0.44, and 0.42 for hair color, tattoo shape, and both, compared to 0.43, 0.59, and 0.70 for age, education, and both. This suggests that LLMs are sensitive to the contextual importance of additional features when determining allocations, meaning that in real applications, reductions in stratification are conditioned on the quality of known features in available data.
Together, these results highlight both the promise and the limitations of structural interventions. Fixing low success rates or introducing job heterogeneity can mitigate stratification with certain prompts, but ideal conditions are only attained when trading-off believability. Adding richer contextual features is more principled, but this is conditioned on the availability of salient features, and some models remain stubbornly anchored to spurious signals even when the most indicative features are provided. Overall, structural modifications provide partial leverage on stratification but do not guarantee robustness.
5.3 Explicit incentivization via prompt steering
Our last series of interventions focuses on prompt steering to reduce stratification. We test four steering prompts targeting different aspects of the LLM’s allocation decisions: directly instructing the model to be fair, emphasizing the LLM’s internal values such as equality and fairness, describing broader societal values of fairness in the city, and adding an explicit diversity term to the objective function. The internal value steer was placed in the system prompt, while the others were added to the user prompt describing the hiring setup. Details on prompts and modifications are in Appendix A.2.
Unlike with prior interventions, the fourth steer (targeting the model’s objectives) was extremely effective across direct and CoT prompts (Figure 7), while also being simple to implement in practice (unlike structural interventions). While Gemini remained biased, remarkably, almost all other models and prompts had SI values lower than both the random baseline and humans fulfilling the same objective. In contrast, the other steering interventions were sometimes successful but did not reduce stratification nearly as much (Figure 7)111Claude 4 Sonnet refused to respond after the internal value steer under both direct and CoT prompts. . This contrast reinforces that while LLMs can align with general value statements, they are far more effective when the incentive of acting in line with such values is concrete and measurable. Our findings return us to the theme of LLMs being great optimizers—demonstrating that as models become better at following instructions to complete tasks, the objectives they follow must evolve with them to achieve desired social outcomes.
6 Discussion
Our results indicate that LLMs demonstrate a new kind of bias —the creation of novel stereotypes—which manifests over repeated interactions in stateful frameworks. Through carefully designed experiments inspired by social science literature, we show how LLMs are even more prone than humans to develop such biases, even when underlying differences do not exist. While much of the fairness literature focuses on measuring inequality through the lens of representational bias (Blodgett et al., 2020), our work demonstrates the consequences of allocational bias, i.e., the unequal distribution of outcomes and opportunities, that can stem from the decisions of large language models, which in turn lead to novel representational distortions that reinforce and legitimize these distributive disparities over time.
Counter to existing literature and bias benchmarks, our results reveal that newer and more capable LLMs stratify more severely than their predecessors in identical sequential decision-making scenarios. One simple reason for this trend is that better models draw more precise inferences about past outcomes. Instead of choosing randomly, a more advanced LLM may favor job candidates from a group when earlier assignments of similar jobs to that group succeeded. However, this reasoning-based tendency can be maladaptive, as it risks reducing exploration and, in turn, inadvertently marginalizing certain social groups. As LLMs become increasingly capable at optimizing toward a given objective, it is essential to define that objective carefully; while AI systems may succeed in domains with clear ground truth, in social domains where truth is often indeterminate, it is more desirable to thoroughly explore candidate options before exploiting a seemingly optimal outcome.
Separately, our findings from Section 4.2 suggest a concerning divergence: while more advanced LLMs consistently improve on existing single-turn bias benchmarks (e.g., Parrish et al., 2022), we find the opposite trend in our tests, indicating that current evaluations on single-turn responses may be too isolated to capture the downstream societal outcomes that these models shape over time. Similar to how algorithms shape societal dynamics through feedback loops (O’Neil, 2016), as AI systems become increasingly agentic, they can also construct feedback loops by learning from the outcomes of their own decisions. This shift underscores the need to evaluate LLMs not only via their immediate answers, but also their long-term influence when deployed in continuous, real-world contexts.
Our interventions in Section 5 represent promising strategies to mitigate biases emerging from such feedback loops, but their utility can be limited by requiring unrealistic changes to the environment (e.g., success rates) or reward function (objective steering). Another assumption we make is that groups have equal success rates across all jobs. However, if unequal success rates exist due to covariates such as education, enforcing diversity can reduce overall success (see Appendix G).
More broadly, LLMs’ tendencies to generalize from examples are what enable superior few-shot learning and a myriad of related capabilities—but this ability to extrapolate patterns is the same capacity that drives premature stratification. This raises a central tension in alignment: How do we limit generalization in sensitive cases without suppressing reasoning as a whole? The challenge ahead is to design interventions that selectively discourage harmful pattern-matching while preserving the constructive forms of abstraction that make LLMs powerful. Finding this balance may be far from straightforward, but will pave the way for equitable and socially beneficial AI systems.
Ethics Statement
Our work focuses on analyzing how LLMs may develop social biases through exploration, bringing awareness to practitioners and developers that this is a grounded concern. We envision our work to hopefully help shape a new generation of safer and more robust AI systems, and thus do not envision any negative ethical implications at this time.
Appendix A Prompting Details
A.1 Prompts for primary experiment
In this section, we provide the prompts for our primary hiring experiment, following stimuli from the psychological studies in Bai et al. (2022a; 2025a). The same prompts are used for the temperature intervention and lower success probability intervention. The preamble is shown once at the start. In each round, the hiring round prompt is shown, the LLM responds, and then the round outcome prompt is shown. The direct and chain-of-thought prompts are appended to the hiring round prompt.
A.2 Prompts for diversity steer intervention
A.2.1 Explicit incentivization prompt
The Hiring Round Prompts and the Round Outcome Prompt in this experiment were identical to those shown in Section A.1. The bonus is calculated by
and adding it to the default success base payment of 1 or 0.
A.2.2 System prompt intervention
We employ the default prompts and reward structure in Appendix A.1 for the rest of this intervention.
A.2.3 Encouraging fair decisions
A.2.4 Promoting shared values
We alter the first line of the default preamble prompt as follows, and keep the rest the same.
A.3 Prompts for eliciting model priors on success probabilities
For independent runs, we query the model what it believes the population success rate for a certain job is, with job order being randomly shuffled in each run. We use these probabilities in the realistic job-wise success probabilities intervention.
A.3.1 Elicitation Results
We analyze the elicited job-wise success probabilities by job quadrant (warmth competence; Fiske et al. (2002)). Success probabilities are reasonably correlated with job quadrant. High warmth and high competence jobs such as doctors have the lowest success rate, low warmth and high competence jobs such as lawyers and managers have the second lowest, followed by high warmth and low competence jobs such as receptionists, and finally low warmth and low competence jobs such as janitors and cashiers.
A.4 Prompts and parameters for refugee resettlement experiment (additional information ablation)
Addons for direct and chain-of-thought prompting were identical to those presented in Section A.1.
Respective populations for each of the demographics examined are < 8000 (Udi) (Federal State Statistics Service (Russia), 2024), < 1100 (Ket) (Vajda, 2007), < 300 (Taz) and < 800 (Tofa) (Federal State Statistics Service (Russia), 2010). All of them are extremely low-population indigenous peoples living in the Siberian and Caucasus regions of Russia and Central Asia, justifying their use as demographic labels that minimize the risk of prior associations or stereotypes carrying over into the experiments.
We ran four different sub-experiments, each taking on a different level of information: demographic only, demographic and age, demographic and education, demographic and education and age. We ran trials of 40-round hiring simulations for each scenario. In each round, the age and education attributes for each candidate were sampled randomly and independently from the attributes listed below, adopted from Bansak et al. (Bansak et al., 2018).
A.5 Prompts and parameters for military conscription assignment experiment
Appendix B Metrics: Analyses and Interpretations
For each metric presented in Section 3.2, we perform controlled and representative numerical experiments to present more tangible interpretations for their respective range of values.
B.1 Stratification Index
B.1.1 Relation to Mutual Information
Under certain conditions, our Stratification Index (SI) is equivalent to mutual information (MI). Specifically, this occurs when job categories occur equally as frequently (assumption 2). We prove this below.
Lemma 1 (Equivalence of SI and MI under uniform job category marginals).
Let be a random variable for demographic group, for job class, and for run of the experiment. Assume that:
1.
Job classes take values in a finite set with .
2.
For each run , the marginal job distribution is uniform on , i.e.
Define the Stratification Index (SI) as
where is the uniform distribution on and is the Shannon entropy (with log base 2), then
i.e., SI equals the expected mutual information between and across runs. In particular, in a single-run (when is constant), we have
Proof.
Fix an arbitrary run . We write all quantities conditioned on and then average over at the end.
First, note that by definition of conditional entropy,
Therefore, for this fixed run ,
Plugging this into the inner expression of equation 4, we obtain
Next, use the uniform-marginal assumption. For each run , we have
so the entropy of the job variable given is
Substituting equation 9 into equation 8 yields
where the last equality is precisely the definition of the conditional mutual information between and given :
Now take expectation over on both sides. Using equation 4 and the above identity, we obtain
In the special case where there is only a single run (or is almost surely constant), conditioning on becomes redundant and the equality reduces to
where we again use the assumption that is uniform, so .
This completes the proof. ∎
B.1.2 Empirical Validation
SI is intended to measure to what degree each demographic is funneled into its own particular set of jobs. Empirically, to measure how well SI adheres to this trend, we design a controlled experiment where in a trial of 40 rounds, each demographic is assigned its main “quadrant” of jobs, where different demographics can be assigned to the same quadrant. Note that this means in some trials, certain quadrants will not be mapped to, and so we do not draw jobs from those quadrants. In each round out of 40, with probability , we select the demographic that maps to that quadrant (if there are multiple, choose from the applicant subset randomly) and with probability , we choose a random demographic.
B.2 Between-Group Divergence
BGD is intended to measure how different the job distributions are across demographics. To measure this, we design a controlled experiment where each demographic is mapped to its own “main” quadrant such that a bijection is formed. For each group’s hires, we form a distribution over quadrants as a mixture between uniform and disjoint allocation:
This means that with all groups have identical uniform distributions, while with each group concentrates entirely on its assigned quadrant. Intermediate values of tilt each group’s distribution toward its own quadrant while retaining some mass elsewhere. A small proportion of hires are then randomly reassigned to add noise. From these distributions, we compute the average Jensen–Shannon distance between groups, which increases as rises, reflecting greater between-group divergence.
B.3 Group Assignment Stochasticity Index
GASI is intended to measure how stable group–quadrant mappings are across repeated runs. In the controlled experiment, each run begins by choosing the mapping rule: with probability we use a fixed universal mapping of groups to quadrants, and with probability we generate a random one-to-one mapping. Within that run, jobs are drawn from the set of occupations in each quadrant, and the group hired is the one assigned to that quadrant under the current mapping. This produces a distribution over quadrants for each group in each run. GASI is then computed as the average Jensen–Shannon distance between distributions of the same group across runs. When , group–quadrant assignments vary randomly across runs, so distributions for a given group differ widely and GASI is high. When , assignments are consistent across runs, so each group’s distribution converges and GASI is low. Thus GASI decreases as increases, capturing the stability of group–quadrant associations.
Appendix C Human Participants
In this section, we describe the demographics of the humans comprising our baseline (originally collected in Bai et al. (2025a)). As stated in their paper, the human data is collected with the following details:
1.
1310 participants were sourced from the CloudResearch High-Quality Subject pool (cloudresearch.com). All speak English as their first language and are at least 18 years old (mean age = 40).
2.
51% of the participants were female, 46% were male, and 1% were non-binary.
3.
74% of participants were White, 10% Black, 6% Hispanic, 5% Asian, and 4% multiracial.
4.
75% of participants hold some college/bachelor degree.
5.
The average political orientation of the participants was 3.94 (1 = extremely conservative, 6 = extremely liberal).
These demographics reflect typical characteristics of online American workers for psychological studies. Crucially, the core result in Bai et al. (2025a) () holds when controlling for individual differences in age, gender, race, education, and political orientation.
Of these 1310 participants, 600 were relevant to our human baselines: 200 for the classic setting, 200 for the altered setting with , and 200 for the diversity steer intervention.
Appendix D Rank-Ordered Allocation Matrices (Hiring Experiment)
In this section, we show how newer-generation models tend to stratify more than older models. We do this for six families of models: Gemini, GPT, Claude, Llama 3.2, Llama 4, and Qwen-2.5. In each rank-ordered allocation matrix, higher stratification is closer to the identity matrix, while lower stratification is closer to uniform spread (see example comparison below).
D.1 Gemini Model Family
Gemini 1.5 Flash Direct
Gemini 1.5 Flash CoT
Gemini 2.0 Flash Direct
Gemini 2.0 Flash CoT
![[Uncaptioned image]](x16.png)
Gemini 2.5 Flash Direct
![[Uncaptioned image]](x17.png)
Gemini 2.5 Flash CoT
![[Uncaptioned image]](x18.png)
D.2 GPT Family
GPT-3.5 Direct
![[Uncaptioned image]](x19.png)
GPT-3.5 CoT
![[Uncaptioned image]](x20.png)
GPT-4o Direct
![[Uncaptioned image]](x21.png)
GPT-4o CoT
![[Uncaptioned image]](x22.png)
D.3 Claude Family
Claude 3 Haiku Direct
![[Uncaptioned image]](x23.png)
Claude 3 Haiku CoT
![[Uncaptioned image]](x24.png)
Claude 4 Sonnet Direct
![[Uncaptioned image]](x25.png)
Claude 4 Sonnet CoT
![[Uncaptioned image]](x26.png)
D.4 Llama 3.2 Family (varying by size)
Llama 3.2 3B Direct
![[Uncaptioned image]](x27.png)
Llama 3.2 3B CoT
![[Uncaptioned image]](x28.png)
Llama 3.2 11B Direct
![[Uncaptioned image]](x29.png)
Llama 3.2 11B CoT
![[Uncaptioned image]](x30.png)
Llama 3.2 90B Direct
![[Uncaptioned image]](x31.png)
Llama 3.2 90B CoT
![[Uncaptioned image]](x32.png)
D.5 Llama 4 Family
Llama 4 Scout Direct
![[Uncaptioned image]](x33.png)
Llama 4 Scout CoT
![[Uncaptioned image]](x34.png)
Llama 4 Maverick Direct
![[Uncaptioned image]](x35.png)
Llama 4 Maverick CoT
![[Uncaptioned image]](x36.png)
D.6 Qwen-2.5 Family (varying by size)
Qwen-2.5 7B Direct
![[Uncaptioned image]](x37.png)
Qwen-2.5 7B CoT
![[Uncaptioned image]](x38.png)
Qwen-2.5 72B Direct
![[Uncaptioned image]](x39.png)
Qwen-2.5 72B CoT
![[Uncaptioned image]](x40.png)
Appendix E Additional experimental scenarios
We examined the default setup as described in Section 3.1 on two other allocative scenarios: refugee resettlement and military conscript assignment, and observe similarly high levels of stratification as LLMs assigned different demographic groups into systematically distinct roles, suggesting that biased structural patterns persist across domains even when contexts and objectives vary.
Appendix F Prior Biased Associations Experiment
In this section, we provide further evidence that LLMs did not possess any prior beliefs around a relation between the artificial demographic names and job quadrants. We run the hiring game setup in Section 3.1 as follows. For each frontier model (except DeepSeek-R1 and OpenAI o3), prompting method (direct or CoT), and job (20 total), we conduct 20 trials each containing only one job vacancy so as to examine the models’ initial perceptions. Afterwards, we combine all job assignments for each model-prompt combination as a single run of assignments, and calculate the SI for this aggregated run. As shown in Table 2, the SI scores for each model-prompt combination are well below the random baseline, strongly suggesting that the models began without any intrinsic or systematic mapping between demographic labels and job quadrants, confirming that any later structure arises from task dynamics rather than pretrained bias.
Claude Sonnet 4 Gemini 2.5 Flash Llama 4 Maverick GPT–4o Qwen 2.5 72B Prompt CoT Direct CoT Direct CoT Direct CoT Direct CoT Direct Global SI 0.081 0.234 0.037 0.036 0.047 0.142 0.059 0.104 0.026 0.190
Appendix G Objective Demographic-Job Mapping Experiment
In this section, we highlight a challenge of implementing the diversity prompt steer approach demonstrated in Section 5.3. One major limitation of the diversity-bonus intervention is its context-dependence, raising the challenge of knowing when it should be deployed. While explicitly rewarding diversity reduces stratification in synthetic environments, when ground-truth demographic–job mappings do exist, blindly applying this guidance can reduce success rates by penalizing correct allocations, as shown in Figure 13. This challenge is especially acute when the underlying scenario is unknown beforehand, making it difficult to determine whether the intervention is appropriate. As such, although the intervention is valuable for probing the mechanisms behind stereotype emergence, it remains limited as a general-purpose solution, with the central problem being not only how to design interventions, but also how to determine where and when they should be applied.
Appendix H Effects of agentic augmentations
To observe the effect at to which stratification is affected in LLMs when they are equipped with external augmentations commonly integrated with agentic systems, we enable GPT-4o and Gemini 2.5 Flash with the ReAct framework from Yao et al. (2023). In lieu of the chain-of-thought or direct prompting prompts as listed in Appendix A.1, we use the template prompt provided in Yao et al. (2023) enabling the LLM with a tool allowing it to assign a certain demographic to the job opening in a certain round, with the resultant observation being either a successful or unsuccessful outcome.
We still observe highly stratified assignments in both models, with resultant SIs of 1.11 and 1.42 for GPT-4o and Gemini 2.5 Flash, respectively, suggesting that the emergence of stratification is not attenuated by agentic scaffolding such as ReAct, but instead persists across reasoning paradigms.
Appendix I Experiment with real-life demographic labels
To test how LLMs’ emergent biases interact with real demographics and existing stereotypes, we run the original setup described in Section 3, changing the demographics to “White", “Black", “Hispanic", “Asian", and the jobs to stereotypically associated categories from He et al. (2019). Other than these changes, we use the same parameters and prompts in Appendix A.1. The specific jobs are as follows:
1.
White-associated (medicine-related): Doctors, Surgeons, Dentists, Pharmacists, Medical Researchers.
2.
Black-associated (sigmatized): Parking Lot Attendants, Janitors, Sewer Cleaners, Security Guards, Street Vendors.
3.
Hispanic-associated (domestic-related): Housekeepers, Landscapers, Construction Workers, Restaurant Cooks, Nannies.
4.
Asian-associated (science/tech-related): Software Engineers, Data Scientists, Hardware Engineers, IT Specialists, Programmers.
We observe similarly high levels of stratification in GPT-4o and Gemini 2.5 Flash. However, that these patterns are less emergent and more driven by pre-existing social priors, with the resulting allocations exhibit substantially lower GASI values as shown in Table 3, suggesting that in this more socially salient setting the models largely reproduce entrenched associations rather than generating new ones.
| Model | Prompting | SI | BGD | GASI |
| GPT-4o | Direct | 1.52 | 0.75 | 0.14 |
| CoT | 1.21 | 0.65 | 0.28 | |
| Gemini 2.5 Flash | Direct | 1.41 | 0.72 | 0.22 |
| CoT | 1.29 | 0.69 | 0.30 |
Appendix J LLMs’ exploration does not uncover differences between success rates
While previous sections show that LLMs develop allocative biases when demographics have equal performance, in this section we illustrate how this persists even when some demographics perform worse than others. We test this using a modified version of the original hiring setup from Section 3.1, where each of the four demographics is most proficient at their own unique job category (with success probability of 0.9), worst in another unique job category (0.75), and moderate (0.8 and 0.85) at the other two categories.
Once the hiring rounds are complete, we ask the model to identify the demographic group that is most likely to succeed at a randomly sampled job from each of the four categories, as shown below. To prevent anchoring effects, we ask these four questions independently of one another, with the same preceding context containing the prompts and responses from the completed hiring rounds.
We perform experiments with GPT-4o and Gemini-2.5-Flash for both direct and chain-of-thought prompting, and we investigate results in both 40-round and 80-round hiring setups. For each possible combination, we conduct 30 trials. For the 40-hiring-round setups, LLMs only identify the best-performing group 26.3% of the time, barely surpassing random chance. It mistakenly identifies the second-best group 29.0% of the time, the third best group 21.5% of the time, and even the worst-fitting group 23.3% of the time (values and 95% CIs shown in Table 4).
| Model | Prompting | Best | Second-Best | Third | Fourth |
| GPT-4o | CoT | ||||
| Direct | |||||
| Gemini 2.5 Flash | CoT | ||||
| Direct |
We also tested LLMs in an 80-round setting, where the models were told that they had 80 hiring rounds instead of 40, giving them more time to explore. However, we did not notice a statistically significant difference in accuracy vs. the 40-round case (26.2%, 30.5%, 24.4%, 18.9%), suggesting the inability of LLMs to appropriately adapt their exploration patterns in settings that allow for increased exploration to attain better long-term rewards (Table 5).
| Model | Prompting | Best | Second-Best | Third | Fourth |
| GPT-4o | CoT | ||||
| Direct | |||||
| Gemini 2.5 Flash | CoT | ||||
| Direct |